Blue spruce Corp. has just purchased equipment that requires annual payments of $55,000 to be paid at the end of each of the next 4 years. The appropriate discount rate is 10%. What is the present value of the payments

A. 174343

B. 220144

C. 107119

D. 210279
Answer #1

The Correct Answer is A. 174343

Calculation of Present Value of Payments

To find the present value we will need to multiply the cash flows of every year with the discounting factor. The total of the present value column will be the present value of the payments.

Present Value
Years Cash flow Discounting factor @ 10% Present Value
1 55000 0.9091 50000
2 55000 0.8264 45455
3 55000 0.7513 41322
4 55000 0.6830 37566
174343

Note - Discounting factor can simply be calculated by the following steps

For year 1 = 1 / 1.10

= 0.9091

For year 2 = 0.9091 / 1.10

= 0.8264

For year 3 = 0.8264 / 1.10

= 0.7513

For year 4 = 0.7513 / 1.10

= 0.6830